### v1

@@ Appendix A Proofs of Main Theorem
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}\geq x\right)\leq\mathbb{P}(|M|>C)+\mathbb{P}\left(|M|<C,|M|\max_{t}\left\|f_{it}\right\|_{2}\cdot\|e_{t,ik}\|_{2}\geq x\right).$
Choosing $x\asymp d_{k}\sqrt{\Theta_{ii}}$, we have with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}-T^{-c_{5}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}\leq C_{6}\cdot\frac{d_{k}\sqrt{\Theta_{ii}}}{T}.$
@@
$\|\Delta_{3}(g_{i\ell},\ell\neq k)\|_{2}\leq$ $\left\|\frac{1}{T}\sum_{t=1}^{T}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}-\mathbb{E}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}\}}\right\|_{2}$
$+\left\|\frac{1}{T}\sum_{t=1}^{T}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}+\left\|\mathbb{E}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}\}}\right\|_{2}$
$\leq$ $C_{7}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\Theta_{ii}}+C_{7}\cdot\frac{d_{k}\log(d)\sqrt{\Theta_{ii}}}{T}.$
@@
$\widehat{\Sigma}_{ik}$ $=\widehat{\Sigma}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
$=$ $\widetilde{\lambda}_{i,i}a_{ik}a_{ik}^{\top}+\frac{1}{T}\sum_{t=1}^{T}f_{it}\left(\prod_{\ell\neq k}^{K}a_{i\ell}^{\top}\widehat{g}_{i\ell}^{(m)\top}\right)a_{ik}\otimes\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}\widehat{g}_{i\ell}^{\top}\right)$
$+\frac{1}{T}\sum_{t=1}^{T}f_{it}\left(\prod_{\ell\neq k}^{K}a_{i\ell}^{\top}\widehat{g}_{i\ell}^{(m)\top}\right)\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}\widehat{g}_{i\ell}^{\top}\right)\otimes a_{ik}+\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes{\cal E}_{t}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
$+\sum_{j\neq i}^{r}\widetilde{\lambda}_{j,i}a_{jk}a_{jk}^{\top}+\sum_{j_{1}\neq j_{2}}^{r}\frac{1}{T}\sum_{t=1}^{T}f_{j_{1}t}f_{j_{2}t}\otimes_{k=1}^{K}a_{j_{1}k}\otimes_{k=K+1}^{2K}a_{j_{2}k}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
$+\frac{1}{T}\sum_{t=1}^{T}\sum_{j\neq i}^{r}f_{jt}\otimes_{k=1}^{K}a_{jk}\otimes{\cal E}_{t}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}+\frac{1}{T}\sum_{t=1}^{T}\sum_{j\neq i}^{r}f_{jt}{\cal E}_{t}\otimes_{k=K+1}^{2K}a_{jk}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
@@
$\left\|\frac{1}{\widetilde{\lambda}_{i,i}}\Psi_{1}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})T}\sum_{t=1}^{T}f_{it}a_{ik}\otimes\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\right\|_{2}\leq C_{4}(\psi^{\text{\footnotesize ideal}})^{2}+C_{4}(\sqrt{r}\psi_{0})\psi^{\text{\footnotesize ideal}},$ (85)
$\left\|\frac{1}{\widetilde{\lambda}_{i,i}}\Psi_{2}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\otimes a_{ik}\right\|_{2}\leq C_{4}(\psi^{\text{\footnotesize ideal}})^{2}+C_{4}(\sqrt{r}\psi_{0})\psi^{\text{\footnotesize ideal}}.$ (86)
@@
$\left\|\frac{1}{\widetilde{\lambda}_{i,i}}a_{ik,\perp}\Psi a_{ik}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}a_{ik,\perp}^{\top}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\otimes a_{ik}\right]a_{ik}\right\|_{2}$
@@
$\sup_{u\perp a_{ik}}\left|u^{\top}\left(\widehat{a}_{ik}\widehat{a}_{ik}^{\top}-a_{ik}a_{ik}^{\top}\right)a_{ik}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\otimes a_{ik}\right]a_{ik}\right|$
@@
In addition, by (A) and (80),
$\sup_{u}\left|u^{\top}P_{a_{ik},\perp}\widehat{a}_{ik}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\otimes a_{ik}\right]a_{ik}\right|$
@@
$\sup_{u}\left|u^{\top}(\widehat{a}_{ik}-a_{ik})-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\right]\right|$
@@
$\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\right]=\frac{1}{\Theta_{ii}}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}b_{i\ell}^{\top}\right)\right].$
@@
$\left|\left(\widehat{a}_{ik}^{\top}a_{ik}\right)^{2}-1-\frac{1}{\Theta_{ii}^{2}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})^{2}}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\right]^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\right]\right|$
@@
The proof of the consistency of $\widehat{r}^{\rm uer}$ draws on methods similar to those in Ahn and Horenstein, (2013) and Han et al., 2022b , given that our CP tensor factor model can be equated to a vector factor model. Moreover, the consistency of $\widehat{r}^{\rm ip}$ aligns with the proofs in Han et al., 2022b , as our CP tensor factor model can also be regarded as a Tucker factor model with a uniform Tucker rank of $(r,...,r)$. Specifically, lemmas akin to Lemmas 11 and 12 (or Lemmas 14 and 15) in Han et al., 2022b can be derived under our assumptions. It leads to $\mathbb{P}(\widehat{r}_{k}=r,1\leq k\leq K)\to 1$. We omit the detailed proofs as they are laborious, albeit straightforward, adaptations for a specialized case of the Tucker factor model. ∎
@@ Appendix B Techinical Lemmas
Let $M\in\mathbb{R}^{d_{1}\times d_{2}}$ be a matrix with $\|M\|_{\rm F}=1$ and ${a}$ and ${b}$ be unit vectors respectively in $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$. Let $\widehat{a}$ be the top left singular vector of $M$. Then,
@@ Appendix C More Simulation Results
In this section we show the simulation results of Configuration I, II and III with AR coefficient on $g_{it}$, $\phi$, equal to $0.5$. We can see that AC-ISO algorithm has a better performance since the signal strength in the auto-covariance grow with $\phi$. CC-ISO algorithm, however, outperforms AC-ISO algorithm even with stronger serial correlation in the factor process.

### v2

@@ Appendix A Proofs of Main Theorem
$\mathbb{P}\left(\left\|\sum_{t=1}^{T}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\geq x\right)$
$\leq$ $\mathbb{P}(|M|>C)+\mathbb{P}\left(|M|<C,|M|\max_{t}\left\|w_{i}f_{it}\right\|_{2}\cdot\|e_{t,ik}\|_{2}\geq x\right).$
Choosing $x\asymp\sqrt{d_{k}r}(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\Theta_{ii}}$, we have with probability at least $1-\exp(-T^{c_{2}})-T^{-c_{2}}-T^{-c_{5}}$,
$\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}\leq C_{6}\cdot\frac{\sqrt{d_{k}r}(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\Theta_{ii}}}{T}.$
@@
$\|\Delta_{3}(g_{i\ell},\ell\neq k)\|_{2}\leq$ $\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}-\mathbb{E}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\leq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$+\left\|\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}+\left\|\mathbb{E}w_{i}f_{i,t}e_{t,ik}^{\top}\mathbf{1}_{\{\|e_{t,ik}\|_{2}\geq C\sqrt{d_{k}}(\log T)^{(\vartheta+2)/\vartheta}\}}\right\|_{2}$
$\leq$ $C_{7}\sqrt{\frac{d_{k}\log(d)}{T}}\sqrt{\Theta_{ii}}+C_{7}\cdot\frac{\sqrt{d_{k}r}\log(d)(\log T)^{1+\frac{2}{\vartheta}+\frac{1}{\gamma_{1}}}\sqrt{\Theta_{ii}}}{T}.$
@@
Let
$\psi_{0}$ $=\max_{1\leq i\leq r}\max_{1\leq k\leq K}\|\widehat{a}_{ik}^{\rm rcpca}\widehat{a}_{ik}^{{\rm rcpca}\top}-a_{ik}a_{ik}^{\top}\|_{2},$
$\psi$ $=\max_{1\leq i\leq r}\max_{1\leq k\leq K}\|\widehat{a}_{ik}^{{\rm iso}}\widehat{a}_{ik}^{{\rm iso}\top}-a_{ik}a_{ik}^{\top}\|_{2}.$
As $1\lesssim\lambda_{r}$, Theorem 4.2 implies that, in an event $\Omega$ with probability at least $1-T^{-C}-d^{-C}$,
$\psi\leq C_{1}\left(\sqrt{\frac{d_{\max}\log d}{\lambda_{r}T}}+\frac{1}{\lambda_{r}}\right).$
Note that
$\widehat{w}_{i}\widehat{f}_{it}-w_{i}f_{it}$ $=\sum_{j=1}^{r}w_{j}f_{jt}\otimes_{k=1}^{K}a_{jk}\times_{k=1}^{K}\widehat{b}_{ik}^{{\rm iso}\top}+{\cal E}_{t}\times_{k=1}^{K}\widehat{b}_{ik}^{{\rm iso}\top}-w_{i}f_{it}$
$=w_{i}f_{it}\left(\prod_{k=1}^{K}a_{ik}^{\top}\widehat{b}_{ik}^{{\rm iso}}-1\right)+\sum_{j\neq i}w_{j}f_{jt}\prod_{k=1}^{K}a_{ik}^{\top}\widehat{b}_{jk}^{{\rm iso}}+{\cal E}_{t}\times_{k=1}^{K}\widehat{b}_{ik}^{{\rm iso}\top}$
$:={\rm I}_{1}+{\rm I}_{2}+{\rm I}_{3}.$
By (62) in the proof of Theorem 4.2, $\max_{1\leq i,j\leq r}\max_{1\leq k\leq K}|a_{ik}^{\top}\widehat{b}_{jk}^{{\rm iso}}-\mathbf{1}_{\{i=j\}}|\lesssim\psi$ in the event $\Omega$. By (A), it follows that ${\rm I}_{1}\leq C_{1}(w_{i}\psi)$, ${\rm I}_{2}\leq C_{1}(\sqrt{\lambda_{1}}r\psi^{K})$ and ${\rm I}_{3}\leq C_{1}$, in the same event $\Omega$. Since $d_{\max}r\lesssim d$, condition (23) leads to $(\lambda_{1}/\lambda_{r})^{1/2}r\psi^{K-1}\lesssim(\lambda_{1}/\lambda_{r})^{1/2}\psi_{0}^{K-1}\lesssim 1$. Then ${\rm I}_{2}\leq C_{2}(\lambda_{r}^{1/2}\psi)$ in the event $\Omega$. Hence, in the event $\Omega$, $w_{i}^{-1}|\widehat{w}_{i}\widehat{f}_{it}-w_{i}f_{it}|\leq C_{0}(\psi+\sigma w_{i}^{-1})$, which completes the proof. ∎
Proof of Theorem 4.4.
@@
[…]
$+\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}\left(\prod_{\ell\neq k}^{K}a_{i\ell}^{\top}\widehat{g}_{i\ell}^{(m)\top}\right)\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}\widehat{g}_{i\ell}^{\top}\right)\otimes a_{ik}+\frac{1}{T}\sum_{t=1}^{T}{\cal E}_{t}\otimes{\cal E}_{t}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
$+\sum_{j\neq i}^{r}\widetilde{\lambda}_{j,i}a_{jk}a_{jk}^{\top}+\sum_{j_{1}\neq j_{2}}^{r}\frac{1}{T}\sum_{t=1}^{T}w_{j_{1}}w_{j_{2}}f_{j_{1}t}f_{j_{2}t}\otimes_{k=1}^{K}a_{j_{1}k}\otimes_{k=K+1}^{2K}a_{j_{2}k}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
$+\frac{1}{T}\sum_{t=1}^{T}\sum_{j\neq i}^{r}w_{j}f_{jt}\otimes_{k=1}^{K}a_{jk}\otimes{\cal E}_{t}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}+\frac{1}{T}\sum_{t=1}^{T}\sum_{j\neq i}^{r}w_{j}f_{jt}{\cal E}_{t}\otimes_{k=K+1}^{2K}a_{jk}\times_{\ell\in[2K]\backslash\{k,K+k\}}\widehat{g}_{i\ell}^{\top}$
@@
[…]
@@
$\left\|\frac{1}{\widetilde{\lambda}_{i,i}}a_{ik,\perp}\Psi a_{ik}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}a_{ik,\perp}^{\top}\left[\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\otimes a_{ik}\right]a_{ik}\right\|_{2}$
@@
$\sup_{u\perp a_{ik}}\left|u^{\top}\left(\widehat{a}_{ik}\widehat{a}_{ik}^{\top}-a_{ik}a_{ik}^{\top}\right)a_{ik}-\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\otimes a_{ik}\right]a_{ik}\right|$
@@
In addition,
$u^{\top}P_{a_{ik},\perp}\widehat{a}_{ik}=u^{\top}P_{a_{ik},\perp}\widehat{a}_{ik}(\widehat{a}_{ik}^{\top}a_{ik})+u^{\top}P_{a_{ik},\perp}\widehat{a}_{ik}(1-\widehat{a}_{ik}^{\top}a_{ik}).$
By (A) and (82),
[…]
@@
[…]
@@
$\frac{1}{\Theta_{ii}\prod_{\ell\neq k}(a_{i\ell}^{\top}g_{il})}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}g_{i\ell}^{\top}\right)\right]=\frac{1}{\Theta_{ii}}u^{\top}P_{a_{ik},\perp}\left[\frac{1}{T}\sum_{t=1}^{T}w_{i}f_{it}\left({\cal E}_{t}\times_{\ell\in[K]\backslash\{k\}}b_{i\ell}^{\top}\right)\right].$
@@
[…]
@@
The proof of the consistency of $\widehat{r}^{\rm uer}$ draws on methods similar to those in Ahn and Horenstein, (2013) and Han et al., 2022b , given that our CP tensor factor model can be equated to a vector factor model. Moreover, the consistency of $\widehat{r}^{\rm ip}$ aligns with the proofs in Han et al., 2022b , as our CP tensor factor model can also be regarded as a Tucker factor model with a uniform Tucker rank of (r,,r)(r,...,r). Specifically, lemmas akin to Lemmas 11 and 12 (or Lemmas 14 and 15) in Han et al., 2022b can be derived under our assumptions. It leads to $\mathbb{P}(\widehat{r}_{k}=r,1\leq k\leq K)\to 1$. We omit the detailed proofs as they are laborious, albeit straightforward, adaptations for a specialized case of the Tucker factor model. ∎
@@ Appendix B Techinical Lemmas
Let $M\in\mathbb{R}^{d_{1}\times d_{2}}$ be a matrix with $\|M\|_{\rm F}=1$ and a{a} and b{b} be unit vectors respectively in $\mathbb{R}^{d_{1}}$ and $\mathbb{R}^{d_{2}}$. Let $\widehat{a}$ be the top left singular vector of $M$. Then,
@@ Appendix C More Simulation Results
  1. 1.
    In this subsection we show the simulation results of Configuration I, II and III with AR coefficient on $g_{it}$, $\phi$, equal to $0.5$. We can see that AC-ISO algorithm has a better performance since the signal strength in the auto-covariance grow with $\phi$. CC-ISO algorithm, however, outperforms AC-ISO algorithm even with stronger serial correlation in the factor process.
@@
  2. 2.
    We conduct the simulation to evaluate the performance of CC-ISO under Pareto-distributed errors. We set $d_{1}=d_{2}$ and let $d_{k}$ vary over $\{20,40,60,80\}$ and T over $\{100,200,300,400,500\}$. The signal strength is set to $(r-i+1)\sqrt{d}$. Each element in ${\cal E}_{t}$ is generated i.i.d. from Pareto distribution with shape parameter $\alpha=2.5$. Figure 18 shows the boxplots of estimation errors for the factor loadings over 200 repetitions. While Pareto-distributed errors lead to a higher occurrence of outliers compared to normally distributed errors whose results are shown in the main text, CC-ISO remains consistent, demonstrating its robustness in heavy-tailed settings.
    Figure 18: Boxplots of estimation errors with pareto errors over 200 repetitions
  3. 3.
    We conduct simulation for correlated factors. Specifically, we generate data following DGP Setting II from the main text, with $\eta=0.1$. In this setting, $f_{it}$ are initially generated independently across $i$. To introduce correlation among the factors, we pre-multiply $f_{t}$ by a Toeplitz matrix with parameter $\zeta$, defined as
    $f_{t}^{(corr)}$ $=\Sigma_{f}f_{t},$
    $\Sigma_{f,ij}$ $=\zeta^{|i-j|},$
    where $\Sigma_{f,ij}$ denotes the $(i,j)$ entry of $\Sigma_{f}$. By varying $\zeta$, we control the degree of correlation among factors. Figure 19 presents the simulation results, which show that our estimator remains consistent in the presence of correlated factors.
    Figure 19: Boxplots of estimation errors with correlated factors
  4. 4.
    We conduct a simulation study to compare the performance of the initial estimator (RC-PCA) and the iterative algorithm (CC-ISO) under both distinct and non-distinct eigenvalues of $\Sigma_{0}$ (as defined in Algorithm 1). The data-generating process (DGP) is set with dimensions $(d_{1},d_{2},T)=(\bar{d},\bar{d},500)$. In the distinct-eigenvalue setting, we set $\eta=0.1$ so that $\delta\leq 0.2$ and $w_{i}=(r-i+1)\sqrt{d}$, where $d=d_{1}d_{2}$. For the non-distinct-eigenvalue setting, following Setting V in the main text, we set $w_{i}=d^{\alpha/2}$ and fix $\delta=0$. Figures 20 and 21 present the results for these two settings, respectively. In both settings, the CC-ISO algorithm outperforms RC-PCA in terms of estimation error and variance. In particular, due to non-zero $\delta$ in the distinct-eigenvalue setting, RC-PCA estimates loadings with a non-vanishing bias, but this bias is removed by CC-ISO.
    Figure 20: Boxplots of estimation errors of RC-PCA and ISO with distinct eigenvalues of $\Sigma_{0}$
    Figure 21: Boxplots of estimation errors of RC-PCA and ISO with non-distinct eigenvalues of $\Sigma_{0}$
  5. 5.
    We conduct simulation study to evaluate Theorem 4.4(i) with estimated variance $\widehat{\sigma}_{u,ik}$ and $\widehat{\Theta}$, which are proposed in Chen et al., (2025), our accompanying paper on the forecast with tensor data. Specifically, we propose a thresholding estimator $\widehat{\Sigma}_{e}^{{\cal T}}$:
    $\widehat{\Sigma}_{e}^{{\cal T}}={\cal T}_{\lambda}\left(\frac{1}{T}\sum_{t=1}^{T}\widehat{e}_{t}\widehat{e}_{t}^{\top}\right),$
    where $\left(\widehat{\Sigma}_{e}^{\cal T}\right)_{(j,l)}={\cal T}_{\lambda}\left(\frac{1}{T}\sum_{t=1}^{T}\widehat{e}_{jt}\widehat{e}_{lt}\right)$, ${\cal T}_{\lambda}(\cdot)$ is a thresholding operator and $\widehat{e}_{t}$ is the vectorized estimated error using the CC-ISO algothrithm. Then we can estimate the variance $\sigma_{u,ik}^{2}$ via $\widehat{\sigma}_{u,ik}^{2}=\widehat{h}_{ik}^{\top}\widehat{\Sigma}_{e}^{{\cal T}}\widehat{h}_{ik}$, where $\widehat{h}_{ik}=\widehat{b}_{iK}\odot\cdots\odot\widehat{b}_{i,k+1}\odot\widehat{P}_{a_{ik},\perp}u\odot\widehat{b}_{i,k-1}\odot\cdots\odot\widehat{b}_{i1}$, and $\widehat{\Theta}=\widehat{W}(\frac{1}{T}\sum_{t=1}^{T}\widehat{F}_{t}\widehat{F}_{t}^{\top})\widehat{W}$. All settings are the same as Configuration (VII) in the main text but we let $T=\lceil 800+d^{3/4}\rceil$ for the consistency of $\widehat{\sigma}_{u,ik}$. The results are shown in Figure 22.
    Figure 22: Histogram of $\sqrt{\widehat{\Theta}_{ii}T}\widehat{\sigma}_{u,ik}^{-1/2}u_{l}^{\top}\left(\widehat{a}_{ik}-\operatorname{sign}\left(\widehat{a}_{ik}^{\top}a_{ik}\right)\cdot a_{ik}\right)$ with weak cross-sectionally correlated errors.
  6. 6.
    We conduct a simulation study to evaluate the performance of the CC-ISO algorithm under unbalanced dimensions. The data is generated according to DGP Setting II in the main text, with $\eta=0.1$. We fix $d_{2}=40$ while varying $d_{1}$ from 20 to 200 and $T$ from 100 to 500. In this setup, since the noise tensor does not exhibit cross-sectional dependence, the rate from Remark 4.3 can be simplified to $\max_{i}\|\widehat{a}_{ik}^{{\rm iso}}\widehat{a}_{ik}^{{\rm iso}\top}-a_{ik}a_{ik}^{\top}\|_{2}=O_{\mathbb{P}}(\sqrt{d_{k}/(dT)})$, with $d=d_{1}d_{2}$. As $T$ increases, the estimation errors for both $\widehat{a}_{i1}(1\leq i\leq r)$ and $\widehat{a}_{i2}(1\leq i\leq r)$ decrease. Furthermore, with $d_{2}$ fixed, an increase in $d_{1}$ improves the convergence rate for the second-mode loading vectors $\widehat{a}_{i2}^{{\rm iso}}$, while the convergence rate for the first-mode loading vectors $\widehat{a}_{i1}^{{\rm iso}}$ remains unchanged. The simulation results, presented in Figure 23, are consistent with this theoretical analysis.
    Figure 23: Boxplots of estimation errors under unbalanced dimensions. The two cases, $k=1$ and $k=1$, represent the estimation error $\max_{i}\|\widehat{a}_{ik}^{{\rm iso}}\widehat{a}_{ik}^{{\rm iso}\top}-a_{ik}a_{ik}^{\top}\|_{2}$ for each respective mode.
  7. 7.
    We measure the computation time of the proposed algorithms, RC-PCA and CC-ISO, and compare them against TPCA Babii et al., (2022) and the GE method proposed by Chang et al., (2023) with $K=3$. Simulations are conducted with data dimensions $(80,80,500)$ and rank $r=3$. To evaluate the computation cost of Randomized Projection, we consider two settings: one with well-separated eigenvalues and another with closely spaced eigenvalues. The tuning parameters are set to $L=2r^{2}$, $c_{0}=0.1$ and $\nu=0.8$. Table 4 reports the median computation time and the median number of iterations (of CC-ISO) over 200 repetitions.
    In the well-separated eigenvalues setting, RC-PCA requires more time than TPCA and GE due to the computation of $\Sigma_{0}$, which has a dimension of $6400\times 6400$ in our DGP. However, this additional cost yields a benefit: RC-PCA attains lower estimation errors, especially when factor loadings are not orthogonal. In the close-eigenvalue setting, the computational cost of RC-PCA increases further because the Randomized Projection step involves performing SVD at least $2\times L$ times. Nevertheless, this added complexity leads to improved estimation performance in difficult scenarios with nearly equal eigenvalues, highlighting a trade-off between computational efficiency and accuracy. For CC-ISO, each iteration takes approximately 2.7 seconds, which is comparable to the computation time of TPCA.
    DGP RC-PCA CC-ISO # Iteration TPCA GE, K=3
    Distinct Eigenvalues 16.94 8.39 3 2.26 6.48
    Close Eigenvalues 131.19 8.19 3 2.13 6.41
    Table 4: Median of computation time (in second) of RC-PCA (column 1), CC-ISO(column 2), Tensor PCA(column 4), and generalized eigen-analysis based estimation (column 5) , along with the median number of iterations (column 3) for CC-ISO. The first row presents results for the distinct-eigenvalue setting. The second row corresponds to the setting with close eigenvalues, where Randomized Projection is applied.
  8. 8.
    We conducted a robustness check for $c_{0}$ and $\nu$. Using Setting V in the paper with $\bar{d}=20$ and $T=500$, we vary $c_{0}\in[0.05,0.5]$ and $\nu\in[0.5,0.9]$. Figure 24 reports the results, with the first box in each group showing the error of the initial estimators without randomized projection (C-PCA). RC-PCA performs consistently across all values of $c_{0}$ and $\nu$, and clearly outperforms C-PCA.
    Figure 24: Boxplots of estimation errors, C-PCA vs. RC-PCA, across different $c_{0}$ and $\nu$.
Appendix D More Empirical Results
Figure 25: The mode-wide eigenvalue ratio plot (the first column) and the mode-wide scree plot (the second column) of the excess return of characteristic decile portfolios, starting from the second largest eigenvalue. The first row shows the plot of the first mode, i.e., the characteristic mode, and the second row shows the second mode, i.e., the decile mode. In the scree plot, the y-axis indicates the proportion of the $i^{th}$ largest eigenvalue to the sum of eigenvalues. In the eigenvalue ratio plot, the y-axis indicates the ratio of the $i^{th}$ eigenvalue to the $(i+1)^{th}$ eigenvalue. The plots suggest the rank pair $(2,3)$ for the Tucker factor model.
Appendix E More Discussion on Initial Factor Estimates
Theorem 4.2 relies on initial factor estimates being reasonably close to the ground truth. This requirement is, in fact, standard and typically unavoidable from a theoretical standpoint. Least-squares estimation for both Tucker and CP factor models involves non-convex optimization problems with many local optima, most of which are not close to the ground truth; see, for example, Chen and Fan, (2023); Han et al., (2024). This challenge is well recognized and is a main reason why estimation of tensor factor models is notoriously difficult.
In the optimization field, there are two general strategies to handle this issue. One approach is to incorporate occasional jumps into the iterations to enable escape from poor local optima (see, e.g., Belkin et al., (2018)). The other is to ensure that the initialization is close to a good local optimum. Our iterative algorithm 3 is related to the second strategy. Similarly, the HOSVD used in Lettau, (2024) provides a high-quality initialization close to the ground truth, which can be refined by the iterative HOOI algorithm. Since HOOI also falls into the second category, its performance crucially depends on good initialization (Zhang and Xia,, 2018). As highlighted in Kolda and Bader, (2009) and other related literature, theoretical convergence guarantees for alternating least squares in CP decomposition are still lacking. In this sense, our RC-PCA (Algorithm 1) plays a role analogous to HOSVD, while our ISO procedure (Algorithm 3) serves as the CP analogue of HOOI in the Tucker factor model analyzed in Lettau, (2024).